billy madison consumes. 100 units of x and 50 units of y. the price price of x roses from 2 to 3. the price of y remains at 4?



a) how much must billy's income rise so that he can exactly afford 100 units of x and 50 units of y




1
Expert's answer
2021-11-15T11:38:44-0500

a)

"income=p_xx+p_yy\\x=100,y=50\\p_x=2,p_y=4\\income=2(100)+4(50)=400"now px rises to 3 , but x=100,y=50

"income=3(100)+4(50)=500\\"

income must rise by "=500-400=100"